OpenStudy (anonymous):

What is the best way to begin solving the following system by substitution? -7x - 2y = -13 x - 2y = 11

OpenStudy (jackellyn):

Solve for variable y in the first of second equation. Lets do the second x-2y=11 -2y = 11-x y = (11-x)/-2 y = -11/2 + x/2 y = 1/2x - 11/2 Substitute into first equation -7x-2(1/2x-11/2)=-13 -7x -x +11 = -13 -8x = -13 -11 -8x = -24 x=3 Which you can substitute into the second eqn (3)-2y=11 -2y = 11-3 -2y = 8 y=-4